OpenStudy (anonymous):

**Algebraic Fractions** Express this as a single fraction in its simplest form 15x^2 / y * y^3 / x (Easier to read if you right it down) Could I have a walkthough? :)

OpenStudy (anonymous):

\[\frac{15x^2}{y}\frac{y^3}{x}\] can be rewritten as: \[\frac{15x^2}{x}\frac{y^3}{y}\] Note that \[\frac{a^m}{a^n}=a^{m-n}\] So we can break this into two parts: \[\frac{15x^2}{x}=15x^{2-1}=15x\] And the second part: \[\frac{y^3}{y}=y^{3-1}=y^2\] Bringing it all together: \[15xy^2\]

OpenStudy (anonymous):

15xy^2 You have ((15x^2)/(y)) * ((y^3)/(x)) Cancel the powers and you're left with 15xy^2

OpenStudy (anonymous):

Thank you! This makes sense. I have some others with brakcets that I may need help on.

OpenStudy (anonymous):

Feel free to post a fresh question when you get stuck on the next one. That way it gets attention from anyone looking :)

OpenStudy (anonymous):

Sorry but I have another one like this but with a / instead of a * can I still rewrite the same?